### v2

# Challenges in numerical integration in physics-informed neural networks modelling

Abstract
In this paper, we numerically examine the challenges that emerge in numerical integration for various tasks now tackled by physics-informed neural networks. Specifically, we illustrate how inaccurately computed integrands can cause serious precision issues. A major difficulty lies in detecting these problems, since a simple large-scale view of the integrand may not reveal any potential errors, and the resulting outcomes are often mistakenly considered correct. To address this, it is critical to determine whether standard double-precision arithmetic suffices for evaluating the integrand or if higher precision is necessary. We evaluate and suggest numerical quadrature methods for selected examples and parameter values, particularly focusing on challenging scenarios.
Received 6 February 2025
Keywords: variable precision arithmetic; numerical integration; adaptive quadrature; physics-informed neural networks
MSC classification: 65D30; 65G30; 68T07
ACM classification: G.1.0; G.1.4; I.2.6; J.2
JEL classification: C63
1 Introduction
Floating-point numbers are widely used in physics-related numerical calculations because they are easy to understand in this context. They have become the standard method for representing real numbers in computers. The IEEE-754 Standard, established in 1985, has significantly contributed to the widespread adoption of floating-point arithmetic by providing clear and practical guidelines for its implementation. To recall the history of floating-point arithmetic as well as its specification mandated by the IEEE-754 Standard we refer to the recent survey by Boldo, Jeannerod, Melquiond, and Muller (2023).
However, in many mathematical models we can observe that IEEE-754 Standard 32-bit or 64-bit floating-point format (denoted binary32 or binary64 respectively) is not always fully sufficient, see for example works by Bailey and Borwein (2005), who studied especially applications in physics (Bailey, Barrio, and Borwein, 2012; Bailey and Borwein, 2015). Computations that require higher than double precision for robust and exact decision making were introduced in Pal, Koul, Musadeekh, Ramakrishna, and Basu (2004). For a brief overview of numerical quadratures and their failures we refer the reader to the Section 4 in Daněk and Pospíšil (2020).
Physics-informed neural networks (PINNs) modelling involves among others a numerical calculation of integrals that usually depend on several model parameters. Optimization tasks consist of large number of integral evaluations with high precision and low computational time requirements. However, for some model parameters, many numerical quadrature algorithms fail to meet these requirements. We can observe an enormous increase in function evaluations, serious precision problems and a significant increase of computational time.
To overcome the problems caused by the floating-point arithmetic limits, high-precision or variable-precision arithmetic is a rapidly growing part of scientific computing environments. In the above mentioned papers we can find among others currently available software packages for high-precision floating-point arithmetic. For example in MATLAB, there exists a possibility of defining variables and perform numerical calculations in variable-precision arithmetic using the vpa bundle that is part of the Symbolic Math Toolbox (according to the documentation, vpa was introduced already before R2006a). Variable in the name suggests that a user can set the number of significant digits arbitrary, by default it is 32 significant digits. Numerical integration using variable precision was introduced later in MATLAB R2016b, however, it can be used only for functions that can be easily converted to symbolic expressions.
In recent years we can identify a systematic research in the area of computer arithmetic that are regularly presented for example at the IEEE International Symposia on Computer Arithmetic (ARITH) or newly at International Conferences for Next Generation Arithmetic (CoNGA). One area of research focus on correctly rounded evaluation of elementary and special functions, see for example a survey by Brisebarre, Hanrot, Muller, and Zimmermann (2024). Some techniques are based on the error free transformations of the elementary floating point operations (additions, subtraction, multiplication and division) such as TwoSum algorithm by Knuth (1997) that requires 6 floating point operations (flop) or TwoProd algorithm by Dekker (1971) that require 17 flop. These two algorithms are for example used in the compensated Horner’s scheme to evaluate accurately the polynomials (Langlois and Louvet, 2007).
In terms of accurate higher precision numerical quadratures, the research is not at all systematic and the majority papers focus only on solutions tailored to specific problems. For papers with a more general impact we refer to the results about Gauss–Legendre quadrature by Fousse (2007a), Newton–Cotes quadratures by Fousse (2007b) or Gauss–Turán quadrature by Gautschi (2014). A comparison of three high-precision quadrature schemes was performed on several testing examples by Bailey, Jeyabalan, and Li (2005), namely the Gaussian quadrature (quadgs), an error function based quadrature scheme (quaderf), and a $\tanh$$\sinh$ quadrature (quadts) were analysed and compared.
In our original conference paper Daněk and Pospíšil (2015) we show that even a very simple quadratic function can produce serious numerical problems if evaluated incorrectly. In a subsequent paper Daněk and Pospíšil (2020) we analysed more complicated integrands that occur in inverse Fourier transforms integrals. Although the motivation comes from the option pricing problems in mathematical finance, we can observe similar misbehaviour in many multi-physics and multi-scale models as well. In this paper we show that in some cases it is actually necessary to use the high or variable precision arithmetic (vpa) in order to get correct integration results that are crucial for the training of neural networks. The aim of this paper is to show how similar problems arise in PINNs modelling and how to handle the precision within desired bounds.
The structure of the paper is as follows. In Section 2, we introduce the inaccurately evaluated functions of interest. Both a rather complicated integrand from the mathematical finance as well as a simple quartic polynomial examples are presented. In Section 3, we compare different numerical quadratures applied to studied integrals and suggest a usage the variable precision arithmetic in problematic cases. In the second part of the paper we study the heat equation with one spatial variable where the initial condition is our inaccurately evaluated function of interest. In Section 4 we show how inaccuracy influences the analytical solution and in Section 5 how it behaves in the numerical approximation of the solution calculated using PINNs. We conclude in Section 6.
2 Inaccurately evaluated functions
[…]
In the past we witnessed already several examples when erroneous navigation input to the rocket’s guidance system (not only in Ariane 5) lead to the placement of satellites to the wrong orbit. Many other particular examples not only from physics can be found for example in the papers by Bailey and Borwein (2005); Bailey, Barrio, and Borwein (2012); Bailey and Borwein (2015).
For the purposes of presented manuscript we analyse two different examples of functions that in the models have to be numerically integrated for which an accurate evaluation is crucial. In Subsection 2.1 we start with a motivating example coming from mathematical finance that was for the first time introduced by Daněk and Pospíšil (2020). To avoid rather complicated formulas in order to clarify the main message of the present manuscript, in Subsection 2.2 we describe in detail how a simple polynomial of fourth order can be inaccurately evaluated very easily.
2.1 Inverse Fourier transform integrand example
Let us now consider a motivating example coming from mathematical finance that was for the first time studied by Daněk and Pospíšil (2020). It can be shown (Baustian, Mrázek, Pospíšil, and Sobotka, 2017) that in the so called approximate fractional stochastic volatility jump-diffusion (AFSVJD) model the price of a European call option with strike price $K$ and time to maturity $\tau$ can be written using the inverse complex Fourier transforms integral as
[…]
with $X=\ln\frac{S}{K}+r\tau$, fundamental transform $\hat{F}(k,v,\tau)=\exp(C(k,\tau)+D(k,\tau)v)$ and
$C(k,\tau)$ $=\kappa\theta\underbrace{\left(Y\tau-\frac{2}{B^{2}}\right)}_{C_{1}(k,\tau)}\underbrace{\ln\left(\frac{1-ge^{-d\tau}}{1-g}\right)}_{C_{2}(k,\tau)},\quad D(k,\tau)=Y\frac{1-e^{-d\tau}}{1-ge^{-d\tau}},\quad Y=-\frac{k^{2}-ik}{b+d},$
$g$ $=\frac{b-d}{b+d},\quad d=\sqrt{b^{2}+B^{2}(k^{2}-ik)},\quad b=\kappa+ik\rho B,\quad B=\varepsilon^{H-1/2}\sigma,$
$\phi(k)$ $=\exp\left\{-i\lambda\beta k\leavevmode\nobreak\ \tau+\lambda\tau\biggl{[}\hat{\varphi}(k)-1]\biggr{]}\right\},\quad\hat{\varphi}(k)=\exp\left\{i\mu_{J}k\leavevmode\nobreak\ -\frac{1}{2}\sigma_{J}^{2}k^{2}\right\},$
$\beta$ $=\exp\{\mu_{J}+\sigma_{J}^{2}/2\}-1.$
In this paper we omit all the mathematical finance details and focus strictly on the integrand $f(k)$ and its numerical integral later in Subsection 3.1.
In Figure 1 we consider market data for European call options to the FTSE 100 index dated 8 January 2014, see (Pospíšil and Sobotka, 2016), namely $\tau=0.120548,K=6250,r=0.009,S=6721.8$ and model parameters as listed directly in the picture. Similarly as Daněk and Pospíšil (2020) we demonstrate the numerical misbehaviour by changing only the values of $\sigma$ and also the zoom depth. Whereas the inaccurately enumerated values are in double precision arithmetic (red), the smooth values are in vpa (blue). In detailed zoom we can see that for greater values of $\sigma$, inaccuracies are more of local character, and for decreasing $\sigma$ we observe not only oscillating imprecision, but also a severe numerical bias, see bottom pictures for $\sigma=0.000001$. Bottom right picture with even more detailed zoom of the double evaluated integrand reveals inaccurate numerical oscillations that are not visible on the less detailed zoom on the left.
Figure 1: Function $f(k)$ for some values of $\sigma$.
2.2 Problematic quartic function example
Let us now consider a simple continuous bounded real function $\phi(x)$ of one real variable
$\phi(x)$ $=(b^{2}(x)-a^{2}(x))/\delta,$ (2)
where
$a(x)$ $=x+\delta^{2},$ (3)
$b(x)$ $=\sqrt{a^{2}(x)+\delta(x-1)^{2}(x+1)^{2}}.$ (4)
Function $\phi(x)$ depends internally on some real parameter $\delta>0$. This particular example has its simplified form
$\phi(x)=(x-1)^{2}(x+1)^{2}$ (5)
that even does not depend on $\delta$. It is worth to mention that such simplification is usually not possible like in the example in Section 2.1. We denote the MATLAB implementations of both forms as phi_double() (that implements (2)) and phi_exact() (that implements (5)), see Figure 2.
function value=phi_double(x,opt)
delta = opt.delta;
a = x + delta.^2;
b = (a.^2+delta.*(x-1).^2.*(x+1).^2).^.5;
value = (b.^2-a.^2)./delta;
end
function value=phi_exact(x,opt)
% delta = opt.delta;
value = (x-1).^2.*(x+1).^2;
end
Figure 2: MATLAB implementation of functions phi_double() and phi_exact().
We further introduce the so called variable precision arithmetic (vpa), also called high or arbitrary precision arithmetic. In MATLAB, vpa(x,d) uses variable precision arithmetic to evaluate each element of the symbolic input x to at least d significant digits, where d is the value of the digits function, whose default value is 32. MATLAB implementation phi_vpa() of our function in variable precision arithmetic is listed in Figure 3. Please note that the high precision value is at the end converted back to double.
function value=phi_vpa(x,opt)
delta = opt.delta;
a = vpa(x + delta.^2);
b = vpa((a.^2+delta.*(x-1).^2.*(x+1).^2).^.5);
value = vpa((b.^2-a.^2)./delta);
value = double(value);
end
Figure 3: MATLAB implementation of function phi_vpa().
Figure 4 demonstrates the inaccuracy in evaluating the function values for different values of the parameter $\delta$. Whereas a global view to the function $\phi(x)$ for values of delta less than ca $10000$ (top left picture) does not show any discrepancies, a detailed zoom (top right picture) indicates the differences in evaluating the function values in double (blue) and vpa (red) arithmetic. It is worth to remind that the differences between exact and vpa (converted back to double) values are within the machine precision. Increasing values of $\delta$ further increases the inaccuracies of the phi_double(). For $\delta$ equal to $225000$ (bottom right picture) or greater phi_double() returns zero for all arguments.
Figure 4: Detailed views to the inaccurately evaluated function phi_double() (blue) in comparison to the accurate phi_vpa() (red) for different values of $\delta$.
3 Numerical quadratures and their failures
The aim of this section is to show how numerical quadratures fail to numerically evaluate the integral of functions introduced in the previous section within a prescribed precision. Moreover, if adaptive quadratures are used, we observe an enormous increase of function evaluations and hence the increase of computational time.
Tested quadratures (existing functions in MATLAB):
integral:
adaptive Gauss quadrature, in particular GaussKronrod(7,15),
quad:
adpative Simpson quadrature,
quadl:
adaptive Lobatto quadrature,
trapz:
trapezoidal numerical integration,
vpaintegral:
adaptive Gauss quadrature, again GaussKronrod(7,15), but in variable precision arithmetic, or more precisely in symbolic operations, i.e. the integrand must be convertible to a symbolic expression.
Adaptive numerical quadratures either fail or in more recent versions of MATLAB at least warn you that the approximate bound on error is huge or that the maximum number of function evaluations was reached. In the tables below we will denote these two types of warnings using numbered superscripts as
Warning1: Reached the limit on the maximum number of intervals in use. Approximate bound on error is $\dots$. The integral may not exist, or it may be difficult to approximate numerically to the requested accuracy.
Warning2: Maximum function count exceeded; singularity likely.
Failure of the adaptiveness of the general Gauss-Kronrod($n,2n+1$) quadrature on even a piece-wise constant function is described in Example 4.1 in Daněk and Pospíšil (2020) where authors suggest how to set the adaptive refinement in order to satisfy the local error estimate to be within the prescribed relative error, but even if this local error bound is zero, the value of the calculated integral does not have to be precise at all.
Further, by integral_vpa we denote the method in which the integrand (values of the integrated function) is evaluated in the variable precision arithmetic with the result being converted back to double. Similarly we denote quadl_vpa and trapz_vpa.
For all quadratures we set the same error tolerances, absolute tolerance AbsTol=1e-12 and relative tolerance RelTol=1e-8 with only one exception, namely for vpaintegral default values AbsTol=1e-10 and RelTol=1e-6 were used.
Following the same structure as in the previous section, in Subsection 3.1 we will first numerically integrate the function $f(k)$ coming from mathematical finance (introduced in Subsection 2.1) and in Subsection 3.2 the numerical integral of function $\phi(x)$ (introduced in Subsection 2.2). Since this function $\phi(x)$ will serve as the main example later in the paper, in Subsection 3.3 we also show how inaccuracy of the quadratures changes when the function is multiplied by the Fourier basis functions.
3.1 Inverse Fourier transform integration
Let us now evaluate the integral in (1), in particular
$I=\int\limits_{0+i/2}^{L+i/2}f(k)\,\mathrm{d}k,$ (6)
where $L=100$ is chosen for demonstration purposes only. Due to the complicated form of function $f(k)$ that depends on many parameters, see (1), the exact result is not known. In Table 1 we can see comparison of different quadratures applied to (6) that is depicted in Figure 1. In the last column, Fevals denote the number of function (integrand) evaluations, i.e. the number of adaptive refinements. As we can see, for all values of $\sigma$, integral_vpa, quadl_vpa and trapz_vpa gives results comparable within the machine precision and computational times. The value of integral_vpa is therefore taken as the reference value and the value Error in the third column is the difference to this value. If vpa is not used, all quadratures are naturally much faster, but less precise. It is worth to emphasize that even if calculating the integrand in vpa is time consuming, the computational time can be comparable to the case when the integrand is calculated in double arithmetic only, which is caused by a massive increase of adaptive refinements caused by the inaccurate evaluation and hence the increase of Fevals that are in the order of hundreds of thousands.
To understand the results of vpaintegral one has to realize that its argument is a function that is inaccurately evaluated in double arithmetic and the function tries to convert it to the symbolic expression that cannot be further simplified. Results obtained by vpaintegral in comparison to integral_vpa are less precise and more time consuming, in particular precision worsens and computational time increases with decreasing $\sigma$. As we can see in Table 1 for $\sigma=0.00001$, the computational time of vpaintegral is more than four times higher than for integral_vpa even for weaker tolerances that are described above.
Table 1: Comparison of different quadratures applied to (6) for some values of parameter $\sigma$.
Method Value Error Time [s] Fevals
case $\sigma=0.001$
integral 0.7768149482534377 0.0000001607693588 2.825 271200
integral_vpa 0.7768147874840788 0.0000000000000000 7.486 390
vpaintegral 0.7768147879892994 0.0000000005052205 0.389 1
quad 0.7769560723152419 0.0001412848311630 0.131 5010
quadl 0.9420883388070267 0.1652735513229479 0.082 2024
quadl_vpa 0.7768147874840784 -0.0000000000000004 6.979 236
trapz-0.01 0.7768149982921251 0.0000002108080462 0.014 1
trapz-0.001 0.7768149922690513 0.0000002047849724 0.025 1
trapz_vpa-0.01 0.7768147874840787 -0.0000000000000001 12.131 1
case $\sigma=0.0001$
integral 0.7768395484341176 0.0000247608438022 6.508 584700
integral_vpa 0.7768147875903154 0.0000000000000000 8.523 390
vpaintegral 0.$7$0.0000001181836677 4.652 1
quad 0.7769532420467165 0.0001384544564010 0.146 5016
quadl 0.9352610191414621 0.1584462315511467 0.102 2030
quadl_vpa 0.7768147875903146 -0.0000000000000008 8.152 236
trapz-0.01 0.7768424929317990 0.0000277053414836 0.010 1
trapz-0.001 0.7768394964992882 0.0000247089089728 0.030 1
trapz_vpa-0.01 0.7768147875903151 -0.0000000000000003 12.740 1
case $\sigma=0.00001$
integral 0.7682354116124637 -0.0085793759884747 1.460 94680
integral_vpa 0.7768147876009384 0.0000000000000000 8.577 390
vpaintegral 0.7768968651904418 0.0000820775895034 36.579 1
quad 0.7682368656720007 -0.0085779219289377 0.129 4392
quadl 0.7681158192130139 -0.0086989683879245 0.113 2024
quadl_vpa 0.$7$0.0000000000000007 8.133 236
trapz-0.01 0.7682814251749993 -0.0085333624259391 0.013 1
trapz-0.001 0.7682369706788630 -0.0085778169220754 0.030 1
trapz_vpa-0.01 0.7768147876009375 -0.0000000000000009 11.891 1
3.2 Integration of the problematic quartic function
The goal is to evaluate
$J=\int\limits_{-1}^{1}{\color[rgb]{0,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.5}\phi(x)}\,\mathrm{d}x.$ (7)
Although the exact result ($16/15\approx 1.066666666666667$) of the simplified phi_exact() is known, many numerical quadratures fail to numerically evaluate the integral of phi_double() within a prescribed precision. In contrast to the previous example, when the further simplification of the integrand was not possible and the integration results were not satisfactory, here the automatic conversion of phi_double() to symbolic expression performed by vpaintegral gives exactly the phi_exact() and hence the exact integration result. Further we compare the integration results of quadratures integral, quad, quadl and trapz applied to integrands phi_double(), phi_vpa() and phi_exact(). In Table 2 we can see the results of numerical integration for different quadratures and different values of $\delta$. As we can see all quadratures applied to phi_vpa() or phi_exact() gives the exact results, what differs is the computational time, integral and quad being the fastest. When the quadratures were applied to phi_double() their results were problematic, worsening with increasing $\delta$. For higher $\delta$, the inaccuracy of the integrand evaluation is higher and hence the Fevals in the adaptive quadrature algorithm. In the last column we indicate the appearance of the above mentioned warnings. It is worth to mention that one has to be careful with low computational times that can sometimes occur for inaccurately evaluated integrand. The case $\delta=250000$ gives a completely wrong result $0.000000000000000$, because the integrand is already completely wrongly evaluated, see also Figure 4. Similarly as in the previous example, the massive increase of adaptive refinements is caused by the inaccurately evaluated integrand which leads to a huge increase of Fevals and consequently also the computational times, see for example the case $\delta=100000$ for quad and quadl.
Table 2: Comparison of different quadratures applied to (7) for some values of parameter $\delta$.
Integral Rule $\delta$ Integrand Rule Value Time [s] Fevals Warning
integral 1000 phi_double 1.066666662383775 0.287 12 No
phi_vpa 1.066666666666666 0.175 1 No
phi_exact 1.066666666666666 0.062 1 No
quad 1000 phi_double 1.066666641747667 0.133 14 No
phi_vpa 1.066666666666667 0.245 14 No
phi_exact 1.066666666666667 0.126 14 No
quadl 1000 phi_double 1.066666543018575 0.065 2 No
phi_vpa 1.066666666666667 0.100 2 No
phi_exact 1.066666666666667 0.061 2 No
trapz 1000 phi_double 1.066666666650391 0.058 1 No
phi_vpa 1.066666666666667 4.953 1 No
phi_exact 1.066666666666667 0.054 1 No
integral 10000 phi_double 1.066667217711548 0.146 11 Yes1
phi_vpa 1.066666666666666 0.230 1 No
phi_exact 1.066666666666666 0.073 1 No
quad 10000 phi_double 1.066721252633333 0.160 22 No
phi_vpa 1.066666666666666 0.237 14 No
phi_exact 1.066666666666667 0.125 14 No
quadl 10000 phi_double 1.066649867548471 0.704 128 No
phi_vpa 1.066666666666667 0.105 2 No
phi_exact 1.066666666666667 0.064 2 No
trapz 10000 phi_double 1.066667380000000 0.054 1 No
phi_vpa 1.066666666666667 5.057 1 No
phi_exact 1.066666666666667 0.056 1 No
integral 100000 phi_double 1.073588382586770 0.189 12 Yes1
phi_vpa 1.066666666666666 0.170 1 No
phi_exact 1.066666666666666 0.072 1 No
quad 100000 phi_double 1.075099781351111 25.521 5006 Yes2
phi_vpa 1.066666666666667 0.245 14 No
phi_exact 1.066666666666667 0.134 14 No
quadl 100000 phi_double 1.075775476723900 11.940 2024 Yes2
phi_vpa 1.066666666666667 0.112 2 No
phi_exact 1.066666666666667 0.071 2 No
trapz 100000 phi_double 1.071808512000000 0.062 1 No
phi_vpa 1.066666666666667 5.591 1 No
phi_exact 1.066666666666667 0.063 1 No
integral 250000 phi_double 0.000000000000000 0.082 1 No
phi_vpa 1.066666666666666 0.224 1 No
phi_exact 1.066666666666666 0.078 1 No
quad 250000 phi_double 0.000000000000000 0.104 4 No
phi_vpa 1.066666666666667 0.248 14 No
phi_exact 1.066666666666667 0.139 14 No
quadl 250000 phi_double 0.000000000000000 0.072 2 No
phi_vpa 1.066666666666667 0.137 2 No
phi_exact 1.066666666666667 0.075 2 No
trapz 250000 phi_double 0.000000000000000 0.061 1 No
phi_vpa 1.066666666666667 5.516 1 No
phi_exact 1.066666666666667 0.088 1 No
3.3 Calculation of Fourier coefficients
Let us now consider basis functions
$g_{n}(x)=\sin\left(\pi(2n-1)(x+1)/2\right),\quad n\in\mathbf{N},$ (8)
that are even, orthonormal at $[-1,1]$ and satisfy $g_{n}(-1)=g_{n}(1)=0$. Alternatively, an equivalent formula
$g_{n}(x)=(-1)^{n+1}\cos\left(\pi(2n-1)x/2\right),\quad n\in\mathbf{N},$
can be considered.
The goal now is to evaluate the Fourier coefficients
$A_{n}=\int\limits_{-1}^{1}{\color[rgb]{0,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.5}\phi(x)}g_{n}(x)\,\mathrm{d}x,\quad n\in\mathbf{N}.$ (9)
In Figure 5 we can see how the integrand changes for different values of $n$. The differences in evaluating the integrand values in double (blue) and vpa (red) arithmetic for $\delta=100000$ are visually obvious.
Figure 5: Graph of inaccurately evaluated integrand $\phi(x)g_{n}(x)$ for $\delta=100000$ and different values of $n$.
In Tables 3 and 4 we list results for $A_{1}$ and $A_{2}$ (9) respectively. We can observe a similar behaviour as in Table 2 where the integrand did not contain the basis function. As we can see, basis functions do not bring additional difficulties. Due to the boundedness of the basis functions, the numerical difference between the double and vpa can be smaller, but the product ${\color[rgb]{0,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.5}\phi(x)}g_{n}(x)$ posses the same bad numerical misbehaviour as the inaccurately evaluated ${\color[rgb]{0,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.5}\phi(x)}$.
Table 3: Comparison of different quadratures used to calculate the value of Fourier coefficient $A_{1}$
Integral Rule $\delta$ Integrand Rule Value Time [s] Fevals Warning
integral 1000 phi_double 0.891088555909676 0.148 11 No
phi_vpa 0.891088548280465 0.181 1 No
phi_exact 0.891088548280465 0.082 1 No
quad 1000 phi_double 0.891088497788002 0.210 26 No
phi_vpa 0.891088530616134 0.400 26 No
phi_exact 0.891088530616134 0.202 26 No
quadl 1000 phi_double 0.891088558410162 0.110 8 No
phi_vpa 0.891088548278784 0.191 8 No
phi_exact 0.891088548278784 0.107 8 No
trapz 1000 phi_double 0.891088548289315 0.061 1 No
phi_vpa 0.891088548280465 5.911 1 No
phi_exact 0.891088548280465 0.089 1 No
integral 10000 phi_double 0.891088948393962 0.170 11 Yes1
phi_vpa 0.891088548280465 0.181 1 No
phi_exact 0.891088548280465 0.080 1 No
quad 10000 phi_double 0.891114412663896 0.195 24 No
phi_vpa 0.891088530616134 0.400 26 No
phi_exact 0.891088530616134 0.202 26 No
quadl 10000 phi_double 0.891069033973833 0.109 8 No
phi_vpa 0.891088548278784 0.195 8 No
phi_exact 0.891088548278784 0.107 8 No
trapz 10000 phi_double 0.891089828004092 0.061 1 No
phi_vpa 0.891088548280465 5.551 1 No
phi_exact 0.891088548280465 0.063 1 No
integral 100000 phi_double 0.909820722614225 0.180 12 Yes1
phi_vpa 0.891088548280465 0.200 1 No
phi_exact 0.891088548280465 0.081 1 No
quad 100000 phi_double 0.924049764163183 29.331 5006 Yes2
phi_vpa 0.891088530616134 0.475 26 No
phi_exact 0.891088530616134 0.231 26 No
quadl 100000 phi_double 0.930107304312503 11.368 2018 Yes2
phi_vpa 0.891088548278784 0.193 8 No
phi_exact 0.891088548278784 0.107 8 No
trapz 100000 phi_double 0.907718939549998 0.061 1 No
phi_vpa 0.891088548280465 5.938 1 No
phi_exact 0.891088548280465 0.104 1 No
integral 250000 phi_double 0.000000000000000 0.089 1 No
phi_vpa 0.891088548280465 0.211 1 No
phi_exact 0.891088548280465 0.099 1 No
quad 250000 phi_double 0.000000000000000 0.106 4 No
phi_vpa 0.891088530616134 0.472 26 No
phi_exact 0.891088530616134 0.212 26 No
quadl 250000 phi_double 0.000000000000000 0.075 2 No
phi_vpa 0.891088548278784 0.192 8 No
phi_exact 0.891088548278784 0.105 8 No
trapz 250000 phi_double 0.000000000000000 0.062 1 No
phi_vpa 0.891088548280465 5.544 1 No
phi_exact 0.891088548280465 0.064 1 No
Table 4: Comparison of different quadratures used to calculate the value of Fourier coefficient $A_{2}$
Integral Rule $\delta$ Integrand Rule Value Time [s] Fevals Warning
integral 1000 phi_double -0.132240670973319 0.178 12 Yes1
phi_vpa -0.132240669593892 0.198 1 No
phi_exact -0.132240669593892 0.080 1 No
quad 1000 phi_double -0.132240607546634 0.225 30 No
phi_vpa -0.132240616544284 0.438 30 No
phi_exact -0.132240616544284 0.225 30 No
quadl 1000 phi_double -0.132240665269117 0.107 8 No
phi_vpa -0.132240669544252 0.185 8 No
phi_exact -0.132240669544252 0.105 8 No
trapz 1000 phi_double -0.132240669210701 0.063 1 No
phi_vpa -0.132240669593892 5.482 1 No
phi_exact -0.132240669593892 0.061 1 No
integral 10000 phi_double -0.132241272405812 0.190 11 Yes1
phi_vpa -0.132240669593892 0.181 1 No
phi_exact -0.132240669593892 0.078 1 No
quad 10000 phi_double -0.132263860323737 0.272 38 No
phi_vpa -0.132240616544284 0.441 30 No
phi_exact -0.132240616544284 0.223 30 No
quadl 10000 phi_double -0.132218822360428 0.109 8 No
phi_vpa -0.132240669544252 0.189 8 No
phi_exact -0.132240669544252 0.105 8 No
trapz 10000 phi_double -0.132242712125416 0.060 1 No
phi_vpa -0.132240669593892 5.652 1 No
phi_exact -0.132240669593892 0.069 1 No
integral 100000 phi_double -0.164105328139715 0.212 12 Yes1
phi_vpa -0.132240669593892 0.193 1 No
phi_exact -0.132240669593892 0.083 1 No
quad 100000 phi_double -0.217677705429780 27.429 5006 Yes2
phi_vpa -0.132240616544284 0.505 30 No
phi_exact -0.132240616544284 0.257 30 No
quadl 100000 phi_double -0.175317505741447 11.336 2012 Yes2
phi_vpa -0.132240669544252 0.189 8 No
phi_exact -0.132240669544252 0.110 8 No
trapz 100000 phi_double -0.162243580664938 0.060 1 No
phi_vpa -0.132240669593892 5.521 1 No
phi_exact -0.132240669593892 0.063 1 No
integral 250000 phi_double 0.000000000000000 0.080 1 No
phi_vpa -0.132240669593892 0.245 1 No
phi_exact -0.132240669593892 0.076 1 No
quad 250000 phi_double 0.000000000000000 0.081 4 No
phi_vpa -0.132240616544284 0.436 30 No
phi_exact -0.132240616544284 0.225 30 No
quadl 250000 phi_double 0.000000000000000 0.073 2 No
phi_vpa -0.132240669544252 0.191 8 No
phi_exact -0.132240669544252 0.108 8 No
trapz 250000 phi_double 0.000000000000000 0.060 1 No
phi_vpa -0.132240669593892 5.608 1 No
phi_exact -0.132240669593892 0.062 1 No
4 Heat equation problem
We consider the heat equation in one spatial variable with Dirichlet boundary conditions:
$\dfrac{\partial u}{\partial t}(t,x)-\alpha\dfrac{\partial^{2}u}{\partial x^{2}}(t,x)$ $=0,\quad t\in(0,1],x\in(-1,1),$ (10)
$u(t,-1)=u(t,1)$ $=0,\quad t\in[0,1],$ (11)
$u(0,x)$ $={\color[rgb]{0,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.5}\phi(x)},\quad x\in[-1,1].$ (12)
Solution by separation of variables with $f_{n}(t)=\mathrm{e}^{-\alpha t((2n-1)\pi/2)^{2}}$, base functions $g_{n}(x)$ given by (8) and Fourier coefficients $A_{n}$ given by (9) takes the form
$u(t,x)$ $=\sum_{n=1}^{+\infty}A_{n}f_{n}(t)g_{n}(x).$ (13)
We denote
$u_{n}(t,x)$ $=A_{n}f_{n}(t)g_{n}(x)$ (14)
and the truncated solution after the $N$-th term by
$u^{N}(t,x)$ $=\sum_{n=1}^{N}u_{n}(t,x).$ (15)
In Figure 6 we can see graphs of the first six $u_{n}(t,x)$, $n=1,\dots,6$, (top) given by (14) and $u^{N}(t,x)$ for $N=50$ (bottom) given by (15) in the case when the initial condition ${\color[rgb]{0,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.5}\phi(x)}$ was taken as the phi_exact() or phi_vpa() (the difference between these two cases is again at the order of machine precision and hence only one figure is provided).
Figure 6: Graphs of the first six $u_{n}(t,x)$, $n=1,\dots,6$, (top) and $u^{N}(t,x)$ for $N=50$ (bottom) in the case when the initial condition ${\color[rgb]{0,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.5}\phi(x)}$ was taken as the phi_vpa().
If ${\color[rgb]{0,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.5}\phi(x)}$ is taken as phi_double(), the graphs of $u_{n}(t,x)$ are visually indistinguishable from the exact version presented in Figure 6. However, the graph of $u^{N}(t,x)$ now differs significantly as we can see in Figure 7. Inaccurately evaluated initial condition ${\color[rgb]{0,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.5}\phi(x)}$ and hence the inaccurately evaluated Fourier coefficients $A_{n}$ causes the serious error at the beginning of the time domain. Figure 7 also shows the influence of the truncation $N$, namely that increasing $N$ does not help to solve the problem caused by the inaccurate evaluation described above.
Figure 7: Graphs of $u^{N}(t,x)$ for $N=50$ (top) and $N=500$ (bottom) in the case when the initial condition ${\color[rgb]{0,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.5}\phi(x)}$ was taken as the phi_double().
5 Physics-informed neural network model
We now train a physics informed neural network (PINN) to numerically compute the solution of the heat equation by using the limited-memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS) algorithm.
  • Training data generated on the go:
    • 25 equally spaced time points to enforce each of the boundary conditions $u(t,-1)=0$ and $u(t,1)=0$.
    • 50 equally spaced spatial points to enforce the initial condition $u(0,x)={\color[rgb]{0,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.5}\phi(x)}$.
    • 10,000 interior points to enforce the output of the network to fulfil the heat equation, QRNG (low discrepancy) Sobol sequence.
  • Deep neural network (DNN) architecture:
    • Multilayer perceptron DNN architecture with 9 fully connect hidden layers with 20 neurons each.
    • Input layer has two input channels corresponding to the inputs $x$ and $t$, output layer has one output channel $u(x,t)$.
    • As activation functions we take hyperbolic tangents (for demonstration purposes).
    • Total 3021 learnable parameters, all of type double.
  • Training options (custom train loop):
    • Train for max 3e3 iterations.
    • Stop training when the norm of the gradients or steps are smaller than 1e-12 (never).
    • Use the default options for the L-BFGS solver state.
    maxIterations = 3e3;
    gradientTolerance = 1e-12;
    stepTolerance = 1e-12;
    solverState = lbfgsState;
  • Model loss function: $\text{loss}=\text{MSE}_{f}+\text{MSE}_{u}$, where
    • $\text{MSE}_{f}=\frac{1}{N_{f}}\sum\limits_{i=1}^{N_{f}}\left|f(t_{f}^{i},x_{f}^{i})\right|^{2}$ and $\text{MSE}_{u}=\frac{1}{N_{u}}\sum\limits_{i=1}^{N_{u}}\left|u(t_{u}^{i},x_{u}^{i})-u^{i}\right|^{2}$.
    • Here, $\{t_{u}^{i},x_{u}^{i}\}_{i=1}^{N_{u}}$ correspond to collocation points on the boundary of the computational domain and account for both boundary and initial condition and $\{t_{f}^{i},x_{f}^{i}\}_{i=1}^{N_{f}}$ are points in the interior of the domain.
    • Calculating $\text{MSE}_{f}$ requires the derivatives $\frac{\partial u}{\partial t},\frac{\partial^{2}u}{\partial x^{2}}$ of the output $u$ of the model – automatic differentiation.
    • For MATLAB implementation see Figure 8.
function [loss,gradients] = modelLoss(net,T,X,T0,X0,U0,opt)
TX = cat(1,X,T);
U = forward(net,TX);
gradientsU = dlgradient(sum(U,"all"),{T,X},EnableHigherDerivatives=true);
Ut = gradientsU{1}; Ux = gradientsU{2};
Uxx = dlgradient(sum(Ux,"all"),X,EnableHigherDerivatives=true);
% Calculate mseF. Enforce heat equation.
f = Ut - (opt.alpha).*Uxx;
zeroTarget = zeros(size(f),"like",f);
mseF = l2loss(f,zeroTarget);
% Calculate mseU. Enforce initial and boundary conditions.
TX0 = cat(1,X0,T0);
U0Pred = forward(net,TX0);
mseU = l2loss(U0Pred,U0);
% Calculated loss to be minimized by combining errors.
loss = mseF + mseU;
% Calculate gradients with respect to the learnable parameters.
gradients = dlgradient(loss,net.Learnables);
end
% MATLAB (since R2021a) automatically optimizes, caches, and reuses the traces
accfun = dlaccelerate(@modelLoss);
Figure 8: MATLAB implementation of the modelLoss() function.
In Figure 9, top four pictures show a very good fit of our DNN model to the (exact) solution, with phi_vpa() being used as the initial condition. Bottom four pictures show the influence of inaccurately evaluated initial condition phi_double(). The solution is highly biased. It is worth to mention that in many real world examples we usually do not have an exact solution to compare our fitted DNN to and hence to realize that there is a huge bias is far from being straightforward.
Figure 9: Cut of the solution at four different timestamps. Top four pictures show the solution when phi_vpa() is used as the initial condition, bottom four are for the phi_double() case.
6 Conclusion
Calculating the value of a function is a basic numerical operation in most numerical algorithms. Often, we incorrectly automatically assume that the value calculated is within machine precision in a given arithmetic, not realizing that this may not be the case. Inaccurate function values can subsequently cause very unreliable inputs to algorithms working with that function. Yet we may not even notice that the obtained results are incorrect.
In this article, we have shown examples of how inaccurate function values can negatively influence the results of numerical integration (quadrature) and subsequently modelling using PINNs.
For calculations, we are interested in the accuracy of the result and the computational time. Requested optimal combination is to have an accurate result in minimum time. We have shown in the examples that a longer computational time does not automatically mean a more accurate result, on the contrary, it can cause problems. On the other hand, a sufficiently accurate result may not always be time consuming. It is worth to mention that one has to be careful with low computational times that can sometimes occur for inaccurately evaluated integrand. Therefore, for each computation, it is important to analyze the entire algorithm including all the data.
In contrast to the results of Daněk and Pospíšil (2020), where authors propose a regime switching algorithm to change the integrand evaluation to vpa only in problematic cases, here the PDE solution method is significantly improved only by evaluating the initial condition in vpa.
In this paper we focused on failures of numerical calculations that are based on using the IEEE-754 Standard floating-point formats. In the last decade we can observe also an increase of usage of the IEEE-1788 Standard interval arithmetic floating-point formats or a completely new generalizing approach using the so called universal number format (unum) described for example in the book by Gustafson (2015). Since these approaches are still not widely used in practice we omitted them in the current study, however, it will be interesting to further research their usage not only in the PINNs modelling context.
Numerical integration challenges, as discussed in this paper, arise due to inaccuracies in evaluating integrands, leading to severe precision problems in many models, and as demonstrated in particular in physics-informed neural networks. These issues stem from floating-point arithmetic limitations and improper numerical quadrature methods, especially when model parameters exacerbate these problems. Similar challenges can occur for example in the peridynamic models, which rely on integral formulations to capture long-range interactions and avoid spatial derivatives. The evaluation of integro-differential operators in peridynamics, as seen in its bond-based models, requires precise numerical integration for accurate representation of material behavior, including crack propagation and heterogeneous material response (Haghighat, Bekar, Madenci, and Juanes, 2021; Ning, Cai, Dong, Liu, and Wang, 2023).
In PINNs applied to peridynamics, numerical integration problems may emerge when enforcing physical constraints, such as integro-differential operators, within the loss function. For example, in peridynamic-informed PINNs, inaccurate integration of forces across horizons or improperly handled kernel functions can degrade model accuracy, particularly under complex boundary or material conditions (Ning, Cai, Dong, Liu, and Wang, 2023; Difonzo, Lopez, and Pellegrino, 2024). These problems echo those highlighted in this paper, where detection of precision loss can be non-trivial without detailed numerical diagnostics. Thus, addressing these integration challenges in peridynamic models often necessitates adopting adaptive quadrature methods or higher-precision arithmetic to maintain robustness and accuracy (Ning, Cai, Dong, Liu, and Wang, 2023).
Acknowledgements
Computational resources were provided by the e-INFRA CZ project (ID:90254), supported by the Ministry of Education, Youth and Sports of the Czech Republic.
We thank our colleagues Kamil Ekštein and Ondřej Pražák from the Department of Computer Science and Engineering for their comments related to the implementation of PINNs.